OpenStudy (anonymous):

Trig. Construct an appropriate triangle to complete the table. (0 <_ theta <_ 90 degrees, 0 <_ theta <_ pi/2) Function Theta (deg.) Theta (radia.) Function value tan ? ? 1/sqrt(3) Would the triangle for this be a (30-60-90) triangle or do I have to create my own triangle?? Please help!

OpenStudy (luigi0210):

Wait, are you saying \( tan \theta=\frac{\sqrt{3}}{3}\)?

OpenStudy (anonymous):

No, I need to figure out the theta in degrees and radians for the tangent function in order to complete the table. The function value 1/sqrt (3) is already given.

OpenStudy (anonymous):

But also need to draw a triangle so Im also wondering if I just draw the 30-60-90 triangle or do I create my own triangle?

OpenStudy (anonymous):

The triangle would be a ( 30-60-90) triangle

OpenStudy (anonymous):

thank you! :) would the degree be 30 degrees and pi/6 for the radians?

OpenStudy (anonymous):

Yes